1a. Because his patient is allergic to mineral oil, a physician has asked you to substitute an equal weight of olive oil for the 10 ml of mineral oil he has in a lotion. How many milliliters of olive oil would be needed? Specific Gravity of mineral oil is 0.85 and Specific Gravity of olive oil is 0.91

Ans:_____mL

1b. Rx: Tannic Acid (20.0g)/ Exsicc. Sodium Sulfite (0.2g)/ Glycerin (25g)/ Yellow Ointment (54.8g)/ Total= 100g
How many mL of glycerin must be used to make 100 g of the above preparation? SG of glycerin is 1.25

Ans: ____mL

1c The official formula for Boric Acid Ointment NF XI is as follows: Boric acid powder (100.0 g) Liquid petrolatum (50g) White ointment (850g) How many mL of liquid petrolatum, light, would be needed to make 120 g of this boric acid ointment? Specific Gravity is 0.81

Ans:___mL

1d. One formula for an O/W ointment base is as follows: Calcium citrate(0.05g)  Sodium alginate(3g) Methylparaben (0.2g) Glycerin (45g) Distilled water qs.ad (1000g). How many mL of glycerin would be needed to prepare 500 gm of the above ointment base? Specific Gravity of glycerin is 1.25

1e. What is the specific gravity of a solution having a volume of 473 mL and weighing 500 gm?
Answer #1

1a. Volume of mineral oil is 10 ml. It’s specific gravity is 0.85.

Its mass is = volume * specific gravity

which is 10 * 0.85 = 8.5 grams

This much olive oil is to be added.

Specific gravity of olive oil is 0.91

volume of olive oil = weight / specific gravity

which is 8.5 / 0.91 = 9.34 ml

Volume of olive oil to be added is 9.34ml.

1b. 100 grams of preparation has 25 grams of glycerin in it.

specific gravity of glycerin is 1.25

volume of glycerin is = weight / specific gravity

which is 25 / 1.25 = 20 ml

Volume of glycerin used to male 100 ml of above preparation is 20ml

1c. Components of boric acid ointment NF XI is

Boric acid power 100 g

liquid petrolatum 50 g

White ointment 850 g

Total weight of ointment is 1000gm.

1000grams boric acid ointment has 50 grams liquid petrolatum

120 grams boric acid ointment has 6 grams liquid petrolatum.

Specific gravity of petrolatum is 0.81

volume of petrolatum is = weight / specific gravity

which is 6 / 0.81 = 7.4 ml approximately

Volume of petrolatum needed to make 120 gm boric acid ointment is 7.4 ml.

1d. Components of O/W ointment base is

Calcium citrate 0.05 gm

sodium alginate 3 gm

Methylparaben 0.2

glycerin 45

distilled water 1000g

total weight of O/W ointment base is = sum of weights of all the components

which is 0.05 + 3 + 0.2 + 45 + 1000 = 1048.25 grams

1048.25 grams ointment base has 45 grams glycerin.

500 grams ointment base has [ (45 * 500) /  1048.25 ] grams glycerin

which is 21.46 grams or 21.5 grams approximately

Specific gravity of glycerin is 1.25

volume of 21.5 grams glycerine = weight / specific gravity

which is 21.5 / 1.25 = 17.2 ml approximately

volume of glycerin required to prepare 500 mg of ointment base = 17.2 ml

1e. Volume is 473 ml. Weight is 500 grams.

specific gravity = weight / volume

which is 500/473 = 1.057 gram / ml

This is a brief explanation. Hope this helps you out. Please feel free to give a like and contact us if you need further detail explanation. Thank you and have a great day.